Suppose that a companies stock is expected to pay a dividend of $0.75 at the end of the year and that the dividend is expected to grow at a rate of 7%. The company’s current beta is 1.9, the current risk-free rate is 2.5% and the market risk premium is 8%. What is the intrinsic value of this stock?

$4.24

$10.71

None of these

$7.50

$7.01
Answer #1

Required rate=risk free rate+Beta*market risk premium

=2.5+(8*1.9)

=17.7%

Intrinsic value=D1/(Required return-Growth rate)

=0.75/(0.177-0.07)

=$7.01(Approx).